# A Stock Price Prediction Approach Based on Time Series Decomposition and Multi-Scale CNN using OHLCT Images

### v1

@@ Abstract
Stock price fluctuations are influenced by a variety of factors, including macroeconomic conditions, government policies and market sentiment, which together make price movements complex and difficult to be predicted. Despite many studies aimed at enhancing stock price prediction models, the challenges such as data noise, model overfitting and lack of interpretability are still encountered. To address these issues and improve prediction accuracy, this paper proposes a novel method, named Sequence-based Multi-scale Fusion Regression Convolutional Neural Network (SMSFR-CNN), for predicting stock price movements in the China A-share market.
Firstly, the historical opening price, highest price, lowest price, closing price, and turnover rate (OHLCT) of stocks are converted into the images, separated by weekends, incorporating specific time information to help the CNN learn the impact of different trading periods. Secondly, traditional CNN methods for stock image recognition often suffer from overfitting and difficulty in finding the correct convergence directions. To mitigate this, long sequences of stock features are decomposed into multiple time periods, and OHLCT images at different time scales are utilized as inputs, thus significantly reducing overfitting. Thirdly, in order to overcome the problem that classification labels lose information about the magnitude of stock price changes, we introduce regression labels so that the model can learn more latent features of stock price fluctuations. Meanwhile, we further integrate time series features into the model.
By utilizing CNN to learn sequential features and combining them with image features, we improve the accuracy of stock trend prediction on the A-share market stock dataset. This approach reduces the search space for image features, stabilizes and accelerates the training process. Extensive comparative experiments on 4,454 A-share stocks show that the proposed model achieves 61.15% for positive predictive value and 63.37% for negative predictive value of the stock price trend over the next 5 days, resulting in a total profit of 165.09%.
@@
{graphicalabstract}
{highlights}
SMSFR-CNN achieves 61.15% PPV and 63.37% NPV on 4,454 A-share stocks
Comprehensive comparison of time series baselines and proposed models for stock prediction
Novel OHLCT image format replaces volume with turnover-rate in stock data representation
Weekend information incorporated as blank time separators in OHLCT images
Decomposed approach tackles overfitting in long sequences of stock image features
@@ 1 Introduction
Furthermore, two new architectures are proposed. The first is a Multi-Scale Residual Convolutional Neural Network (MSR-CNN) designed to address the overfitting problem in long sequence images. The second architecture is a Sequence-based Multi-scale Fusion Regression Convolutional Neural Network (SMSFR-CNN), which addresses the problem that using only image features makes it difficult to learn information about stock price fluctuations. By integrating sequence data, SMSFR-CNN better captures stock price trends in the A-share market. We observed that traditional CNN methods often exhibit local feature overfitting and convergence issues in stock image feature extraction, with prediction performance gradually deteriorating as the time of sequence feature maps increases. This observation aligns with the findings of [29]. To address these mentioned issues, we decomposed long sequences of stock image features into multiple time periods according to different time scales and assign different feature weights based on their importance (with higher weights for features closer to the current trading day). This significantly reduces overfitting and enables CNNs to learn long sequence image features better.
@@ 5 Conclusions
A key contribution of this study is the integration of regression labels into the convolutional neural network framework, further proposing the Sequence Multi-Scale Fusion Regression CNN (SMSFR). This approach better learns and predicts the magnitude of stock price changes, which is particularly important in financial markets where the magnitude of price fluctuations is often more critical than the direction. By combining image feature extraction with sequential data, our model can comprehensively understand the complex relationships between different time periods and trading patterns, achieving optimal prediction results.

### v2

@@ Abstract
Recently, deep learning in stock prediction has become an important branch. Image-based methods show potential by capturing complex visual patterns and spatial correlations, offering advantages in interpretability over time series models. However, image-based approaches are more prone to overfitting, hindering robust predictive performance. To improve accuracy, this paper proposes a novel method, named Sequence-based Multi-scale Fusion Regression Convolutional Neural Network (SMSFR-CNN), for predicting stock price movements in the China A-share market.
Firstly, the historical opening, highest, lowest, closing price, and turnover rate (OHLCT) of stocks are converted into images, separated by weekends, with time information to help the CNN learn the impact of different trading periods. To reduce overfitting, long sequences of stock features are decomposed into multiple time periods, and OHLCT images at different time scales are utilized as inputs, significantly reducing overfitting. Thirdly, in order to overcome the problem that classification labels lose information about the magnitude of stock price changes, we introduce regression labels to help the model capture more latent features of stock price fluctuations.
By utilizing CNN to learn sequential features and combining them with image features, we improve the accuracy of stock trend prediction on the A-share market stock dataset. This approach reduces the search space for image features, stabilizes and accelerates the training process. Extensive comparative experiments on 4,454 A-share stocks show that the model achieves 61.15% positive predictive value and 63.37% negative predictive value for the next 5 days, resulting in a total profit of 165.09%.
@@ 1 Introduction
Furthermore, two new architectures are proposed. The first is a Multi-Scale Residual Convolutional Neural Network (MSR-CNN) designed to address the overfitting problem in long sequence images.
The second architecture is a Sequence-based Multi-scale Fusion Regression Convolutional Neural Network (SMSFR-CNN), which addresses the problem that using only image features makes it difficult to learn information about stock price fluctuations. By integrating sequence data, SMSFR-CNN better captures stock price trends in the A-share market. We observed that traditional CNN methods often exhibit local feature overfitting and convergence issues in stock image feature extraction, with prediction performance gradually deteriorating as the time of sequence feature maps increases. This observation aligns with the findings of [29]. To address these mentioned issues, we decomposed long sequences of stock image features into multiple time periods according to different time scales and assign different feature weights based on their importance (with higher weights for features closer to the current trading day). This significantly reduces overfitting and enables CNNs to learn long sequence image features better.
@@ 5 Conclusions
A key contribution of this study is the integration of regression labels into the convolutional neural network framework, further proposing the Sequence Multi-Scale Fusion Regression CNN (SMSFR). This approach can learns and predicts the magnitude of stock price changes better, which is particularly important in financial markets where the magnitude of price fluctuations is often more critical than the direction. By combining image feature extraction with sequential data, our model can comprehensively understand the complex relationships between different time periods and trading patterns, achieving optimal prediction results.